OpenStudy (anonymous):

In y=mx+ b form Is slope -2/3; through (3,2) a)y= -2/3x+ 4 b)y=-2/3x-4

OpenStudy (aroub):

y=mx+b m = -2/3 , (3,2) 2=-2/3(3)+b 2=-2+b 4=b => y=-2/3x+4

OpenStudy (aroub):

(a)